On January 1, 2019, Schooner Corp. issued bonds having the following characteristics: Face value: Issue price: Stated rate: Market rate: Maturity: $400,000.00 $372,431.15 8% per year 9% per year 11 years The bonds make semiannual interest payments on June 30 and December 31 Assume that the company uses the straight-line interest method. Part (a) Prepare the journal entry for interest on June 30, 2019 When computing amounts, use Excel formulas and cell references only. Do not manually input values.) Account Title Debit Credit Part (b) What is the balance of the discount account as of January 1, 2021 (after the fourth interest payment)? (Use Excel formulas and cell references only. Do not manually input values.)
Answer #1
(a) Discount on issue of bonds=Face value-issue price=400000-372431.15=$ 27568.85
Under stright-line interest method, discount to be amortized equally over the life of bond
The bonds make semi-annual interest payments.
Maturity of the bond=11 years
Number of semi-annual interest payments=11*2= 22
Discount to be amorized In a semi-annual interest period=27568.85/22=$ 1253.13
Date Account title Debit Credit
2019
June 30. Interest expense 17253.13
Discount on bonds payable 1253.13
Cash (400000*8%*6/12) 16000
(Journal etry for interest)
(b) Bond issued on Jan 1,2019
Number of semi-annual periods till Jan 1, 2021=4 semi-annual payments
Discount to be amortized for 4 semi-annual payments=4*1253.13=$ 5012.52
Balance of discount account=27568.85-5012.52=$ 22556.33